# Characterising 1-Rectifiable metric spaces via Connected Tangent Spaces

### v1

@@ 5. Connected Tangents Implies Rectifiable
By the virtue of the fact that $Q_{*}$ are nested, $\mathcal{Q}_{*}(\delta,R,C)\subset\mathcal{Q}_{*}(\delta,R+1,C)$, we can assume that $R$ is such that $128/(R+1)<\hat{d}/$ 12. We may then find a further subset $G\subset F$ with
@@
P3.$\quad\sum_{i}\pi_{i}>\frac{1-\varepsilon}{3(2R+3)}\mathcal{H}^{1}(G).$
For a moment fix a $P=(G_{1},G_{2},\hat{\pi},p_{1},p_{2})$ in $\mathcal{{P}}$ and set $\pi=2\hat{\pi}$. Because $\pi<r_{1}$, we can find $(\nu,\rho,y)\in\textnormal{Tan}(\mathcal{H}^{1}\llcorner F,p_{1})$ such that (5.3) holds for $r=\pi$. Then by Theorem 2.26, there exists compact $p_{1}\in K_{F}\subset F\cap B(p_{1},\pi(R+1))$ and $y\in K_{\nu}\subset\operatorname{supp}\nu\cap B(y,R+1)$ such that
M1. $\max\left\{\frac{\mathcal{H}^{1}(F\cap B(p_{1},\pi(R+1))\backslash K_{F})}{\mathcal{H}^{1}(F\cap B(p_{1},\pi))},\nu(B(y,R+1)\backslash K_{\nu})\right\}<\frac{1}{R+1},$
M2. $d_{pmGH}\left((K_{F},\frac{\mathcal{H}^{1}\llcorner K_{F}}{\mathcal{H}^{1}(F\cap B(p_{1},\pi))},p_{1}),(K_{\nu},\nu\llcorner K_{\nu},y)\right)<\frac{3}{R+1}.$
@@
(5.4) $F_{z}\left(\frac{\mathcal{H}^{1}\llcorner\bar{K}_{F}}{\mathcal{H}^{1}(F\cap B(p_{1},\pi))},\bar{\nu}\llcorner\bar{K}_{\nu}\right)+d_{H(z)}(\bar{K}_{F},\bar{K}_{\nu})<\frac{3}{R+1},$
@@
Firstly, note that by our choice of $R$ we now satisfy the assumptions of Proposition 3.4 where $\varepsilon=3/(R+1)$ and $\delta=1/6$, thus there exists $s\in K_{F}\cap B_{d/\pi}(p_{2},1/6)$. This point combined with (5.4), gives us the existence of $\bar{t}\in\bar{K}_{\nu}$ such that $\lambda(\bar{s},\bar{t})\leq 3/(R+1)$. Thus, using the triangle inequality we find that
@@
(5.6) $\frac{\mathcal{H}^{1}\llcorner\bar{K}_{F}(B(\bar{q},1/12))}{\mathcal{H}^{1}(F\cap B(p_{1},\pi))}\geq\nu(B(\bar{q},1/12)\cap K_{\nu})-\frac{3}{R+1}\geq\frac{\hat{d}}{12}-\frac{4}{R+1}.$
@@
$\mathcal{H}^{1}(F\cap B_{d/\pi}(p_{1},R/2)\backslash G_{1}\cup G_{2})$ $\geq(\frac{\hat{d}}{12}-\frac{4}{R+1})\mathcal{H}^{1}(F\cap B(p_{1},\pi))$
$\geq(\frac{\hat{d}}{12}-\frac{4}{R+1})\hat{d}\pi>\frac{\hat{d}^{2}}{24}\pi.$
Since the Besicovitch partitions are disjoint (condition P1.), the balls $B_{d/\pi}(p_{1},R/2)$ corresponding to each partition are disjoint. we may repeat the same process for each partition $P_{i}\in\mathcal{P}$ and use P3. to obtain
$\mathcal{H}^{1}(F\backslash G)$ $\geq\frac{\hat{d}^{2}}{24}\cdot\frac{1-\varepsilon}{3(2R+3)}\mathcal{H}^{1}(G)$
$=\frac{\hat{d^{2}}(1-\varepsilon)}{72(2R+3)}(\mathcal{H}^{1}(F)-\mathcal{H}^{1}(F\backslash G)).$
@@
Finally, we restate the necessary part of [2, Theorem 1.2], with which we obtain our final statement. For $K\geq 1$, let $\textnormal{biLip}(K)^{*}$ be the set of all pointed metric measure spaces whose metric is $K$-bi-Lipschitz equivalent to the Euclidean norm.

### v2

@@ 5. Connected Tangents Implies Rectifiable
By the virtue of the fact that $Q_{*}$ are nested, $\mathcal{Q}_{*}(\delta,R,C)\subset\mathcal{Q}_{*}(\delta,R+1,C)$, we can assume that $R$ is such that $8/(R+1)<\hat{d}/12$. We may then find a further subset $G\subset F$ with
@@
[…]
For a moment fix a $P=(G_{1},G_{2},\hat{\omega},p_{1},p_{2})$ in $\mathcal{{P}}$ and set $\omega=2\hat{\omega}$. Because $\omega<r_{1}$, we can find $(\nu,\rho,y)\in\textnormal{Tan}(\mathcal{H}^{1}\llcorner F,p_{1})$ such that (5.4) holds for $r=\omega$. Then by Theorem 2.29, there exists compact $p_{1}\in K_{F}\subset F\cap B(p_{1},\omega(R+1))$ and $y\in K_{\nu}\subset\operatorname{supp}\nu\cap B(y,R+1)$ such that
M1. $\max\left\{\frac{\mathcal{H}^{1}(F\cap B(p_{1},\omega(R+1))\backslash K_{F})}{\mathcal{H}^{1}(F\cap B(p_{1},\omega))},\nu(B(y,R+1)\backslash K_{\nu})\right\}<\frac{1}{R+1},$
M2. $d_{pmGH}\left((K_{F},\frac{\mathcal{H}^{1}\llcorner K_{F}}{\mathcal{H}^{1}(F\cap B(p_{1},\omega))},p_{1}),(K_{\nu},\nu\llcorner K_{\nu},y)\right)<\frac{3}{R+1}.$
@@
(5.5) $F_{z}\left(\frac{\mathcal{H}^{1}\llcorner\bar{K}_{F}}{\mathcal{H}^{1}(F\cap B(p_{1},\omega))},\bar{\nu}\llcorner\bar{K}_{\nu}\right)+d_{H(z)}(\bar{K}_{F},\bar{K}_{\nu})<\frac{3}{R+1},$
@@
Firstly, note that by our choice of $R$ we now satisfy the assumptions of Proposition 3.5 where $\varepsilon=3/(R+1)$ and $\delta=1/6$, thus there exists $s\in K_{F}\cap B_{d/\omega}(p_{2},1/6)$. This point combined with (5.5), gives us the existence of $\bar{t}\in\bar{K}_{\nu}$ such that $\lambda(\bar{s},\bar{t})\leq 3/(R+1)$. Thus, using the triangle inequality we find that
@@
(5.7) $\frac{\mathcal{H}^{1}\llcorner\bar{K}_{F}(B(\bar{q},1/12))}{\mathcal{H}^{1}(F\cap B(p_{1},\omega))}\geq\nu(B(\bar{q},1/12)\cap K_{\nu})-\frac{3}{R+1}\geq\frac{\hat{d}}{12}-\frac{4}{R+1}.$
@@
$\mathcal{H}^{1}(F\cap B_{d/\omega}(p_{1},R/2)\backslash G_{1}\cup G_{2})$ $\geq(\frac{\hat{d}}{12}-\frac{4}{R+1})\mathcal{H}^{1}(F\cap B(p_{1},\omega))$
$\geq(\frac{\hat{d}}{12}-\frac{4}{R+1})\hat{d}\omega>\frac{\hat{d}^{2}}{24}\omega.$
Since the Besicovitch partitions are disjoint (condition P1.), the balls $B_{d/\omega}(p_{1},R/2)$ corresponding to each partition are disjoint. We may repeat the same process for each partition $P_{i}\in\mathcal{P}$ and use P3. to obtain
[…]
@@
Finally, we restate the necessary part of [3, Theorem 1.2], with which we obtain our final statement. For $K\geq 1$, let $\textnormal{biLip}(K)^{*}$ be the set of all pointed metric measure spaces whose metric is $K$-bi-Lipschitz equivalent to the Euclidean norm.